# mode = title

### v1

@@ Abstract
Endoscopic surgery relies on two-dimensional views, posing challenges for surgeons in depth perception and instrument manipulation. While Simultaneous Localization and Mapping (SLAM) has emerged as a promising solution to address these limitations, its implementation in endoscopic procedures presents significant challenges due to hardware limitations, such as the use of a monocular camera and the absence of odometry sensors. This study presents a robust deep learning-based SLAM approach that combines state-of-the-art and newly developed models. It consists of three main parts: the Monocular Pose Estimation Module that introduces a novel unsupervised method based on the CycleGAN architecture, the Monocular Depth Estimation Module that leverages the novel Zoe architecture, and the 3D Reconstruction Module which uses information from the previous models to create a coherent surgical map. The procedure’s performance was rigorously evaluated using three publicly available datasets (Hamlyn, EndoSLAM, and SCARED) and benchmarked against two state-of-the-art methods, EndoSFMLearner and EndoDepth. The integration of Zoe in the MDEM demonstrated superior performance compared to state-of-the-art depth estimation algorithms in endoscopy, whereas the novel approach in the MPEM exhibited competitive performance and the lowest inference time. The results showcase the robustness of our approach in laparoscopy, gastroscopy, and colonoscopy, three different scenarios in endoscopic surgery. The proposed SLAM approach has the potential to improve the accuracy and efficiency of endoscopic procedures by providing surgeons with enhanced depth perception and 3D reconstruction capabilities.
@@ 1 Introduction
Laparoscopy has become the preferred surgical method for various procedures across medical specialties [1]. However, a significant limitation of this technique is its reliance on two-dimensional views [2], and inferring depth from 2D images can be challenging for surgeons and requires years of experience [3].
@@
To address this challenge, there has been growing interest in integrating 3D imaging technology into surgical procedures. Studies have shown that 3D imaging reduces strain on surgeons, enhances procedural performance, and increases the accuracy of instrument manipulation [4, 5]. These benefits lead to better procedural outcomes and shortened learning curves for surgical trainees [6, 7, 8].
@@
In recent years, camera-based tracking and mapping methods have gained popularity in surgical 3D imaging research [1]. Approaches such as mosaicking, Structure from Motion, and Shape from Template have been explored, but they face limitations in the surgical context. Mosaicking stitches together overlapping images to create a comprehensive view of the scene [9, 10, 11]: however, it struggles with illumination variations, camera motion, moving objects, and noise, which can lead to artifacts and reduced visual quality in surgical environments. Structure from Motion reconstructs 3D structures from 2D images taken from different viewpoints [12, 13]: it faces challenges such as computational complexity, sensitivity to noise and outliers, limited camera calibration, and difficulties handling complex surgical scenes with moving objects, occlusions, and reflections. Shape from Template estimates the 3D shape of a deformable object by aligning a known template to observed images [14, 15], but, it is affected by template size and complexity, image quality and noise, rotation and scaling issues, partial occlusion, and the presence of similar shapes or clutter in surgical scenes.
@@
In contrast, Simultaneous Localization and Mapping (SLAM) has emerged as a promising solution to address these limitations [16, 17, 18]. SLAM algorithms simultaneously estimate the camera pose and construct a map of the environment in real-time, so that they can effectively handle the dynamic and complex nature of surgical scenes, including moving objects, occlusions, and deformations. However, implementing SLAM in endoscopic procedures presents significant challenges. Hardware limitations, such as the use of a monocular camera due to size constraints and the absence of odometry sensors, led SLAM relying solely on visual information for its core processes, such as localization and mapping. Various environmental factors further complicate the application of SLAM, including low-texture surfaces [19], lighting variability [20, 21], organ deformation [22], and the presence of blood, fluids, and smoke [23, 24].
@@
To overcome these limitations in this paper we propose a robust deep learning SLAM approach that effectively works in diverse surgical settings. Our framework combines state-of-the-art and newly developed models, consisting of three main modules:
  1. 1.
    The Monocular Depth Estimation Module (MDEM), which employs a novel deep learning technique that demonstrates exceptional generalization abilities across various imaging contexts.
  2. 2.
    The Monocular Pose Estimation Module (MPEM), using a novel unsupervised method we developed that learns to estimate the relative camera pose between consecutive frames.
  3. 3.
    The 3D Reconstruction Module (3DM), that introduces a multi-phase process that generates point clouds, refines pose estimates, and converts the aligned point clouds into a volumetric representation. The module addresses incremental drift error and scale ambiguity inherent in monocular pose estimation.
To rigorously test our approach and demonstrate its generalization capabilities, we conducted comprehensive evaluations in three surgical environments: laparoscopy, gastroscopy, and colonoscopy. To this end, we used the Hamlyn [17], EndoSlam [16], and SCARED [25] datasets, each offering unique features for validating our proposed method. We also benchmarked our approach against two state-of-the-art methods [17, 16], which are presented in section 5.
@@
The remainder of this paper is organized as follows: section 2 provides an overview of the related work. Section 3 describes our proposed approach, detailing the three primary modules of our framework. Section 4 presents the materials used in this study, including the training, validation, and testing datasets. Section 5 outlines the experimental configuration, including the comparative analysis against the two state-of-the-art approaches, performance metrics, and statistical analysis. Section 6 presents the experimental results. Finally, section 7 provides concluding remarks.
@@ 2.1 Traditional Monocular Visual SLAM Architecture
A classical Monocular Visual SLAM (MVSLAM) system typically consists of several integrated components, as illustrated in Figure 1A.
@@
The framework begins with a feature extraction module, which identifies salient features (like ORB, SURF, and SIFT [26, 27, 28]) within the images. A feature-matching algorithm then establishes correspondences between these features across consecutive frames. The matched features are passed into a motion estimation algorithm and a depth map estimation algorithm to estimate the relative motion between consecutive frames and generate a sparse depth map, respectively. These outputs are utilized by a map-building algorithm to reconstruct the environmental model. Additionally, the MVSLAM framework incorporates an optimization module, which executes procedures such as bundle adjustment [29] and pose graph optimization to refine the map and pose estimates.
@@ 2.2 Challenges and Limitations of Traditional MVSLAM in Endoscopic Surgery
Traditional feature-based methodologies in MVSLAM have encountered notable challenges in the context of endoscopic surgery due to the unique nature of surgical images. Endoscopic images often lack discernible features and are subject to lighting variations, posing substantial difficulties for feature extraction and matching algorithms. For example, a modified version of ORB-SLAM [30], which utilized a limited number of keyframes for dense reconstruction, struggled with soft, texture-less tissue, a common characteristic in Minimally Invasive Surgery (MIS). This highlights the inherent challenge of detecting robust features in environments where texture and contrast are minimal.
Early SLAM techniques, such as those relying on Structure from Motion [31], aimed to concurrently track camera motion and reconstruct the 3D environment. However, these methods were severely limited by the low-texture environments typical of MIS. Structure from Motion relies on robust wide-baseline feature matching, which becomes untenable in texture-deficient surgical scenes. Moreover, the computational expense of Structure from Motion is another drawback, as it requires processing all images collectively to optimize the 3D structures and camera poses, which is not always feasible in real-time surgical applications.
2.3 Deep Learning-based MVSLAM Architecture
@@
2.4 Advancements in Deep Learning-based MVSLAM for Endoscopic Surgery
Early applications of deep learning in MVSLAM for endoscopic surgery have shown promising results. For example, Chen et al. [32] used Generative Adversarial Networks (GANs) for depth estimation, demonstrating the ability to overcome challenges related to feature scarcity, tissue homogeneity, surface deformation, and highly variable specular appearances.
Hybrid approaches, such as the Endo-Depth-and-Motion framework [17], have attempted to integrate deep learning with traditional techniques. This framework employs MonoDepthV2 [33] for depth estimation and a photometric approach inspired by PTAM [34] for pose estimation. While this approach showed improvements over traditional methods, it still encountered limitations in prolonged tasks, as pose estimation was prone to failure over time.
Fully deep learning-based SLAM systems, such as EndoSLAM [16] for gastroscopy, have further advanced the field by employing deep learning methods for both depth and pose estimation. These approaches have demonstrated the potential of deep learning to provide more robust and adaptable solutions for MVSLAM in the surgical environment.
@@ 2.5 Challenges in Deep Learning-based MVSLAM for Endoscopic Surgery
Despite the advancements in deep learning-based MVSLAM for endoscopic surgery, several challenges remain. One major challenge is the need for large amounts of diverse training data to develop models that can generalize well across different surgical scenarios. Collecting and annotating such datasets can be time-consuming and expensive.
Another challenge is the computational complexity of deep learning models, which can hinder real-time performance in resource-constrained surgical environments. Balancing model accuracy and efficiency is crucial for the practical deployment of deep learning-based MVSLAM systems in endoscopic surgery.
2.6 Distinctive Elements of the Proposed MVSLAM Framework
Our proposed MVSLAM framework introduces several distinctive elements that set it apart from the existing literature and address the challenges mentioned above:
  1. 1.
    Integration of the state-of-the-art Zoe model [35] for monocular depth estimation. This model has demonstrated superior generalization capabilities across diverse datasets, enabling our framework to perform well in various surgical scenarios without the need for extensive fine-tuning.
  2. 2.
    A novel pose estimation module, named CyclePose, which leverages the power of Generative Adversarial Networks (GANs) to estimate the relative camera pose between consecutive frames. CyclePose is specifically designed to enhance the accuracy and robustness of pose estimation in the challenging surgical setting, addressing the limitations of previous approaches.
  3. 3.
    A comprehensive 3D reconstruction module that integrates the outputs from the depth and pose estimation modules to generate detailed and coherent 3D models of the surgical scene. The module incorporates several optimization techniques, such as pose graph optimization and the Iterative Closest Point (ICP) algorithm, to refine the pose estimates and enhance the alignment of the reconstructed point clouds. This ensures the generation of accurate and visually appealing 3D models, which can be valuable for surgical planning, guidance, and post-operative analysis.
By introducing these distinctive elements, our proposed MVSLAM framework aims to address the limitations of current methods in endoscopic surgery, such as the inability to handle feature-poor imagery, occlusions by surgical instruments, uniform textures, and variable lighting conditions.
@@ 3 Methods
The proposed MVSLAM approach is composed of three key modules: the Monocular Depth Estimation Module (MDEM), the Monocular Pose Estimation Module (MPEM), and the 3D Reconstruction Module (3DM), as shown in Figure 2.
Figure 2: The approach takes RGB frames as input and outputs a 3D reconstruction of the surgical scene. The MDEM estimates depth maps from RGB frames, while the MPEM, utilizing CyclePose, computes relative motion between consecutive frames, outputting a motion matrix $M=[R,t_{unscaled},1,0]$, where $R$ is the rotation matrix and $t_{unscaled}$ is the unscaled translation vector. The Translation Estimation Module estimates a scaled translation vector $t_{scaled}$, which is then combined with $t_{unscaled}$ using an Unscented Kalman Filter (UKF) to correct the scale of the motion matrix. Finally, the 3DM combines the RGB frames, depth maps, and scaled pose matrices to generate and update the 3D model.
The structure of the framework is designed to ensure that while the MDEM and MPEM operate independently, their outputs synergistically contribute to the effective functioning of the 3DM. The workflow begins with the MPEM (section 3.1), where the RGB input is processed using CyclePose to estimate the relative motion of the camera between consecutive frames, outputting a motion matrix $M=[R,t_{unscaled},1,0]$, where $R$ is the rotation matrix and $t_{unscaled}$ is the unscaled translation vector due to the scale ambiguity inherent in monocular pose estimation. Simultaneously, the MDEM (section 3.2) estimates the depth map of the scene from the RGB input. Given that the translation component ($t_{unscaled}$) of the relative motion matrix is unscaled, a dedicated module, comprising a translation estimation module and an Unscented Kalman Filter, is tasked with scale correction. The translation estimation module estimates a scaled translation vector $t_{scaled}$, which is then combined with $t_{unscaled}$ using the UKF to correct the scale of the motion matrix. The 3DM (section 3.3) then employs the outputs from both the MDEM and MPEM. It utilizes the depth map from the MDEM and the pose estimations from the MPEM to generate an initial point cloud of the current field of view of the endoscopic camera. This point cloud is carefully merged with the broader scene point cloud, leveraging the pose information to ensure precise alignment and integration.
3.1 Monocular Pose Estimation Module (MPEM)
@@
The first step of the approach is based on CyclePose, a modified version of CycleGAN, inspired by RetinaGAN [36] and InfoGAN [37]. The main idea is to learn the estimation of the relative camera pose between two consecutive frames, denoted as $f_{i-1}$ and $f_{i}$, through the pre-task of image-to-image translation in a totally unsupervised manner. We consider this pre-task as a function $G:f_{i-1}\rightarrow f_{i}$, which generates the current frame $f_{i}$ from the previous frame $f_{i-1}$. Similarly, we define the function $F:f_{i}\rightarrow f_{i-1}$, which generates the previous frame $f_{i-1}$ from the current frame $f_{i}$. To improve the performance of this image-to-image translation pre-task, we draw inspiration from InfoGAN’s manipulation of the latent space through the maximization of mutual information. We propose to manipulate the latent space $z$ produced by the generator encoder $E$ by concatenating a prediction of the relative pose $M$ between the two consecutive frames to $z$. The underlying concept is that if the predicted pose $M$ is correct, the image-to-image translation task $G(f_{i-1},M)\approx f_{i}$ will also be successful. Conversely, if the estimated pose is incorrect, the pre-task will fail. To accommodate this, the architecture has been modified at the bottleneck level of the generator architecture (right panel of Figure 3). The modified generator now performs two tasks:
@@
  • Generating the next frame (image-to-image translation pre-task). First, the previous frame $f_{i-1}$ is passed through the encoder $E$. The resulting latent space $z=E(f_{i-1})$ is then concatenated with the predicted pose $M$ from the first task. This concatenated latent space $[z,M]$ is then fed into the generator $G$, which produces the next frame $\hat{F}_{i}=G(z,M)$.
@@
  • The concatenated frame $f_{c}=[F_{i-1},f_{i}]$ is passed through the network to predict the relative pose $M=P(E(F_{c}))$.
  • The previous frame $f_{i-1}$ is then passed through the encoder $E$ for the image-to-image translation task, producing the latent space $z=E(F_{i-1})$.
@@
3.2 Monocular Depth Estimation Module (MDEM)

### v2

@@ Abstract
Endoscopic surgery relies on two-dimensional views, posing challenges for surgeons in depth perception and instrument manipulation. While Monocular Visual Simultaneous Localization and Mapping (MVSLAM) has emerged as a promising solution, its implementation in endoscopic procedures faces significant challenges due to hardware limitations, such as the use of a monocular camera and the absence of odometry sensors. This study presents BodySLAM, a robust deep learning-based MVSLAM approach that addresses these challenges through three key components: CycleVO, a novel unsupervised monocular pose estimation module; the integration of the state-of-the-art Zoe architecture for monocular depth estimation; and a 3D reconstruction module creating a coherent surgical map. The approach is rigorously evaluated using three publicly available datasets (Hamlyn, EndoSLAM, and SCARED) spanning laparoscopy, gastroscopy, and colonoscopy scenarios, and benchmarked against four state-of-the-art methods. Results demonstrate that CycleVO exhibited competitive performance with the lowest inference time among pose estimation methods, while maintaining robust generalization capabilities, whereas Zoe significantly outperformed existing algorithms for depth estimation in endoscopy. BodySLAM’s strong performance across diverse endoscopic scenarios demonstrates its potential as a viable MVSLAM solution for endoscopic applications.
@@ 1 Introduction
Laparoscopy has become the preferred surgical method for various procedures across medical specialties due to its minimally invasive nature [1]. However, a significant limitation of this technique is its reliance on two-dimensional views [2], and inferring depth from 2D images can be challenging for surgeons and requires years of experience [3].
@@
To address this challenge, there has been growing interest in integrating 3D imaging technology into surgical procedures. Studies have shown that 3D imaging reduces cognitive strain on surgeons, enhances procedural performance, and increases the accuracy of instrument manipulation [4, 5]. These benefits lead to better procedural outcomes and shortened learning curves for surgical trainees [6, 7, 8].
@@
In recent years, camera-based tracking and mapping methods have gained popularity in surgical 3D imaging research [1]. Approaches such as mosaicking and Shape from Template have been explored, but each faces specific limitations in the surgical context. Mosaicking, which stitches together overlapping images to create a comprehensive view of the scene [9, 10, 11], struggles with illumination variations, camera motion, moving objects, and noise, which can lead to artifacts and reduced visual quality in surgical environments. Shape from Template estimates the 3D shape of a deformable object by aligning a known template to observed images [12, 13], but, it is affected by template size and complexity, image quality and noise, rotation and scaling issues, partial occlusion, and the presence of similar shapes or clutter in surgical scenes.
In contrast, Simultaneous Localization and Mapping (SLAM) has emerged as a promising solution to address these limitations.
@@
SLAM algorithms simultaneously estimate the camera pose and construct a map of the environment in real-time, so that they can effectively handle the dynamic and complex nature of surgical scenes, including moving objects, occlusions, and deformations. However, implementing SLAM in endoscopic procedures presents significant challenges due to hardware limitations and environmental factors. The use of monocular cameras, necessitated by size constraints, and the absence of odometry sensors force SLAM to rely solely on visual information for localization and mapping. This approach is known as Monocular Visual SLAM (MVSLAM). This reliance on visual data is further complicated by the unique characteristics of surgical images, which often lack discernible feature [14] and exhibit lighting variations [15, 16]. These factors make robust feature extraction and matching particularly difficult for traditional feature-based MVSLAM methods, such as ORB-SLAM [17] and Structure from Motion [18]. Additional complications arise from organ deformation [19], and the presence of blood, fluids, and smoke [20, 21]. These environmental challenges significantly impair the performance of traditional MVSLAM approaches in endoscopic settings.
Recent works [22, 23, 24, 25] have explored deep learning-based approaches to address these challenges. However, these methods, while promising, still struggle with the same environmental factors that plague traditional approaches. Furthermore, these approaches face additional challenges in generalizing to unseen surgical environments due to data scarcity and the wide variability of surgical scenes.
@@
To overcome these limitations in this paper we propose a robust deep learning MVSLAM approach that effectively works in diverse surgical settings, referred to as BodySLAM in the following. It introduces a newly developed model and state-of-the-art that, together, overcome the issues of low-texture surfaces, lighting variations, and generalization to unseen datasets. Our key contributions are:
  • CycleVO: a novel GAN-based pose estimation module that enhances accuracy and robustness in surgical settings while addressing data scarcity. This unsupervised method uniquely applies concepts from CycleGAN and InfoGAN to learn robust pose representations from unlabelled endoscopic video sequences.
  • Integration of the Zoe model [26] for monocular depth estimation: we leverage its superior generalization capabilities across diverse datasets without fine-tuning, marking the first use of such a well-generalized model for depth prediction in endoscopic MVSLAM. We present an in-depth validation procedure demonstrating its robust performance across various endoscopic scenarios.
To rigorously test our approach and demonstrate its generalization capabilities, we conducted comprehensive evaluations using the largest publicly available endoscopic MVSLAM dataset to date. Our experiments span three surgical environments: laparoscopy, gastroscopy, and colonoscopy. We utilized the Hamlyn [22], EndoSLAM [23], and
@@
SCARED [27] datasets, each offering unique features for validating our proposed method. We also benchmarked our approach against four state-of-the-art methods [22, 23, 24, 25], which are overviewed in section 2.3 and then presented in section 5. The remainder of this paper is organized as follows: section 2 provides an overview of the related work. Section 3 describes our proposed approach, detailing the three primary modules of our framework. Section 4 presents the materials used in this study, including the training, validation, and testing datasets. Section 5 outlines the experimental configuration, including the comparative analysis against the two state-of-the-art approaches, performance metrics, and statistical analysis. Section 6 presents the experimental results. Finally, section 7 provides concluding remarks.
@@ 2.1 Traditional Monocular Visual SLAM Architecture
A classical Monocular Visual SLAM (MVSLAM) system typically consists of several integrated components, as illustrated in Figure 1A. The framework begins with a feature extraction module, which identifies salient features (like ORB, SURF, and SIFT [28, 29, 30]) within the images. A feature-matching algorithm then establishes correspondences between these features across consecutive frames. The matched features are passed into a motion estimation algorithm and a depth map estimation algorithm to estimate the relative motion between consecutive frames and generate a sparse depth map, respectively. These outputs are utilized by a map-building algorithm to reconstruct the environmental model. Additionally, the MVSLAM framework incorporates an optimization module, which executes procedures such as bundle adjustment [31] and pose graph optimization to refine the map and pose estimates.
@@ 2.2 Challenges and Limitations of Traditional MVSLAM in Endoscopic Surgery
Traditional feature-based methodologies in MVSLAM have encountered notable challenges in the context of endoscopic surgery due to the unique nature of surgical images. Endoscopic images often lack discernible features and are subject to lighting variations, posing substantial difficulties for feature extraction and matching algorithms. For example, a modified version of ORB-SLAM [32], which utilized a limited number of keyframes for dense reconstruction, struggled with soft, texture-less tissue, a common characteristic in minimally invasive surgery. This highlights the inherent challenge of detecting robust features in environments where texture and contrast are minimal.
Early SLAM techniques, such as those relying on Structure from Motion [33], aimed to concurrently track camera motion and reconstruct the 3D environment. However, these methods were severely limited by the low-texture environments typical of minimal invasive surgert. Structure from Motion relies on robust wide-baseline feature matching, which becomes untenable in texture-deficient surgical scenes. Moreover, the computational expense of Structure from Motion is another drawback, as it requires processing all images collectively to optimize the 3D structures and camera poses, which is not always feasible in real-time surgical applications.
2.3 Recent Advancements in Deep Learning-based Endoscopic MVSLAM
@@
Recent years have witnessed significant advancements in deep learning-based MVSLAM for endoscopic surgery. EndoSLAM [23] represents a fully end-to-end approach, leveraging convolutional neural networks and attention mechanisms for monocular depth and pose estimation in gastroscopy. This work demonstrated the feasibility of deep learning solutions for domain-specific endoscopic SLAM. Hybrid approaches have also emerged, combining deep learning with traditional techniques. Recasens et al. [22] proposed the Endo-Depth-and-Motion framework, which integrates MonoDepthV2 [34] for depth estimation with a PTAM-inspired [35] photometric method for pose estimation. While this approach showed improvements over purely traditional methods, it still encountered limitations in extended procedures. Addressing a critical challenge in endoscopic environments, Shao et al. [24] introduced the concept of appearance flow to handle brightness inconsistency. Their self-supervised framework employs a generalized dynamic image constraint, relaxing the conventional brightness constancy assumption and enhancing robustness to varying illumination conditions. Most recently, Teufel et al. [25] presented OneSLAM, a generalized monocular SLAM approach for endoscopy utilizing tracking any point (TAP) foundation models. This method achieves sparse correspondences across multiple frames for joint optimization of camera poses and anatomical geometry. Notably, OneSLAM demonstrates zero-shot generalization across various endoscopic domains without retraining, while maintaining or surpassing the performance of domain-specific approaches. These advancements illustrate the growing capabilities of deep learning in addressing the unique challenges of endoscopic SLAM, from domain-specific solutions to more generalized approaches that promise broader applicability in surgical environments.
@@ 2.4 Challenges in Deep Learning-based MVSLAM for Endoscopic Surgery
Despite recent advancements, deep learning-based MVSLAM for endoscopic surgery faces several challenges. The complex surgical environment presents unique obstacles that must be overcome [19, 20, 21]. A major hurdle is the requirement for large, diverse datasets to train generalizable models, which are costly and time-consuming to collect and annotate. Additionally, the computational demands of deep learning models can impede real-time performance in resource-limited surgical settings. Striking a balance between model accuracy and efficiency remains crucial for practical implementation.
@@ 3 Methods
Figure 2: The approach takes RGB frames as input and outputs a 3D reconstruction of the surgical scene. The MPEM, utilizing CycleVO, computes relative motion between consecutive frames, outputting a motion matrix $M=[R,t_{unscaled},1,0]$, where $R$ is the rotation matrix and $t_{unscaled}$ is the unscaled translation vector. The MDEM estimates depth maps from RGB frames. The Translation Estimation Module estimates a scaled translation vector $t_{scaled}$, which is then combined with $t_{unscaled}$ using an Unscented Kalman Filter to correct the scale of the motion matrix. Finally, the 3DM combines the RGB frames, depth maps, and scaled pose matrices to generate and update the 3D model
The proposed MVSLAM approach is composed of three key modules (Figure 2): the Monocular Pose Estimation Module (MPEM), the Monocular Depth Estimation Module (MDEM), and the 3D Reconstruction Module (3DM). The structure of the framework is designed to ensure that, while the MPEM and the MDEM operate independently, their outputs synergistically contribute to the effective functioning of the 3DM. The MPEM (section 3.1) processes the RGB input using our novel unsupervised deep network, named as CycleVO, to estimate the relative motion of the camera between consecutive frames, outputting a motion matrix $M=[R,t_{unscaled},1,0]$, where $R$ is the rotation matrix and $t_{unscaled}$ is the unscaled translation vector due to the scale ambiguity inherent in monocular pose estimation. Simultaneously, the MDEM (section 3.2) estimates the depth map of the scene from the RGB input. Given that the translation component ($t_{unscaled}$) of the relative motion matrix is unscaled, a dedicated module, comprising a translation estimation module and an Unscented Kalman Filter, is tasked with scale correction. The translation estimation module estimates a scaled translation vector $t_{scaled}$, which is then combined with $t_{unscaled}$ using the UKF to correct the scale of the motion matrix. The 3DM (section 3.3) then employs the outputs from both the MDEM and MPEM. It utilizes the depth map from the MDEM and the pose estimations from the MPEM to generate an initial point cloud of the current field of view of the endoscopic camera. This point cloud is carefully merged with the broader scene point cloud, leveraging the pose information to ensure precise alignment and integration.
3.1 Monocular Pose Estimation Module (MPEM): CycleVO
@@
The MPEM introduces CycleVO, a novel unsupervised method designed to address the challenge of data scarcity in endoscopic environments. While CycleVO incorporates architectural ideas insipired by CycleGAN [36] for cyclic image translation and concepts from InfoGAN [37] for latent space manipulation, it uniquely applies and extends these concepts to pose estimation. The key innovation lies in learning the estimation of relative camera pose between consecutive frames through the pre-task of image-to-image translation. We define this pre-task as a function $G:f_{i-1}\rightarrow f_{i}$, which generates the current frame $f_{i}$ from the previous frame $f_{i-1}$. Similarly, we define the function $F:f_{i}\rightarrow f_{i-1}$, which generates the previous frame $f_{i-1}$ from the current frame $f_{i}$. To improve the performance of this image-to-image translation pre-task, we draw inspiration from InfoGAN’s manipulation of the latent space through the maximization of mutual information. We propose to manipulate the latent space $z$ produced by the generator encoder $E$ by concatenating a prediction of the relative pose $M$ between the two consecutive frames to $z$. The underlying concept is that if the predicted pose $M$ is correct, the image-to-image translation task $G(f_{i-1},M)\approx f_{i}$ will also be successful. Conversely, if the estimated pose is incorrect, the pre-task will fail. This novel approach creates a synergy between pose estimation and image generation, enabling the model to learn robust pose representations from unlabeled endoscopic video sequences. The CycleVO generator performs two key tasks (right panel of Figure 3):
@@
  • Generating the next frame (image-to-image translation pre-task). First, the previous frame $f_{i-1}$ is passed through the encoder $E$. The resulting latent space $z=E(f_{i-1})$ is then concatenated with the predicted pose $M$ from the first task. This concatenated latent space $[z,M]$ is then fed into the generator $G$, which produces the next frame $\hat{F}i=G(z,M)$.
@@
  • The concatenated frame $f_{c}=[F{i-1},f_{i}]$ is passed through the network to predict the relative pose $M=P(E(F_{c}))$.
  • The previous frame $f{i-1}$ is then passed through the encoder $E$ for the image-to-image translation task, producing the latent space $z=E(F{i-1})$.
@@
3.2 Monocular Depth Estimation Module (MDEM): Zoe